Jim had 11 coins in dimes and quarters. Their value was $1.70.How many of each did he have?
Answer by checkley77(12844)

D+Q=11 OR D=11-Q
.25Q+.10D=1.70
.25Q+.10(11-Q)=1.70
.25Q+1.1-.10Q=1.70
.15Q=1.70-1.10
.15Q=.60
Q=.60/.15
Q=4 NUMBER OF QUARTERS.
D=11-4
D=7 NUMBERS OF DIMES.
PROOF:
.25*4+.10*7=1.70
1.00+70=1.70
1.70=1.70